OpenStudy (anonymous):

A rectangle has an area of 14 square centimeters. Its length is five cm greater than its width. Find the dimensions. A. 2 and 7 B. 4 and 9 C. 6 and 8 D. 7 and 7

OpenStudy (anonymous):

So you can assume that: width = x and length = x+5 So your area = x*(x+5) so you get 14 = x(x+5) Move everything to one side to get x^2 + 5x - 14 = 0 Solve for the roots (x -2)(x + 7) = 0 So x = 2 OR x = -7 But we obviously cannot have a negative dimention. Therefore x = 2 So our width = 2 And our length = x+5 = 7

OpenStudy (anonymous):

So 2 and 7?